The Length of a rectangle is 1 ft more than its width and 1 ft less than the length of a diagonal. Find the length of the diagonal in simplest radical form.
Answer by edjones(8007)

We have a right triangle with the diagonal the hypotenuse.
a^2+b^2=c^2
Let the length be "a".
a=b+1
c=a+1=b+2
(b+1)^2+b^2=(b+2)^2
b^2+2b+1+b^2=b^2+4b+4
2b^2+2b+1=b^2+4b+4
b^2-2b-3=0 Subtract the right side from the left.
(b-3)(b+1)=0
b=3
b+2=5' Ans.